OpenStudy (amtran_bus):

If y<0 and 4x>y, which of the following could be equal to x/y?

OpenStudy (amtran_bus):

A) 4 B) 0 C) 1/4

OpenStudy (agent0smith):

4x>y Divide both sides by 4. Then divide both sides by y

OpenStudy (amtran_bus):

Ok. So x>y/4

OpenStudy (amtran_bus):

Can you show me that last part? I know y is neg so we flip sign

OpenStudy (agent0smith):

\[\large x > \frac{ y }{ 4 }\] divide both sides by y, flip sign \[\large \frac{ x }{ y } < \frac{ 1 }{ 4 }\]

OpenStudy (amtran_bus):

Nice. So it has to be 0